OpenStudy (anonymous):

Visitors to a carnival can buy an unlimited-ride pass for $50 or an entrance-only pass for $20. In one day, 282 passes were sold for a total of $10,680. How many unlimited-ride passes were sold? (2 points)

OpenStudy (anonymous):

Have you been provided with any further information?

OpenStudy (anonymous):

YES GENERALLY YOU R GIVEN MORE INFORMATION

OpenStudy (aripotta):

$50 unlimited-ride passes are going to be variable a $20 entrance-only passes are going to be variable b 282 passes were sold, so a + b = 282 $10,680 worth of passes were sold, so 50a + 20b = 10680

OpenStudy (aripotta):

this is a system of equations problem. our equations being a + b = 282 50a + 20b = 10680

OpenStudy (aripotta):

use the first equation and solve for b

OpenStudy (anonymous):

Solve @AriPotta

OpenStudy (aripotta):

once you do that, you can substitute that for b in the second equation and solve for a

OpenStudy (aripotta):

a + b = 282 b = 282 - a 50a + 20b = 10680 50a + 20(282 - a) = 10680 solve for a and you'll have your answer

OpenStudy (anonymous):

you decided to use subsitution method

OpenStudy (aripotta):

yes, that seemed like the best way to go, in my opinion

OpenStudy (anonymous):

I would have gone for multiplying by -20

OpenStudy (aripotta):

well, it doesn't really matter. either way, you'll get your answer :P

OpenStudy (anonymous):

lets go then

OpenStudy (aripotta):

i want @tqosborne to get the answer for him/her self

OpenStudy (anonymous):

-20 a -20b = -20(282) 50a + 20b = 10, 680

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

its up to you @tqosborne

OpenStudy (anonymous):

@AriPotta leave this one to you!

OpenStudy (anonymous):

Sorry I was busy doing another problem. Thanks for the help guys :)